# Branching Brownian motion versus Random Energy Model in the supercritical phase: overlap distribution and temperature susceptibility

### v2

@@ 3.4 The REM case
proving that the function $Q(\,\cdot\,,\beta^{\prime})$ is almost surely right differentiable at $1$. On the other hand, Theorem 1.1 implies that $F(\,\cdot\,,\beta^{\prime})=\mathbb{E}[Q(\,\cdot\,,\beta^{\prime})]$ has an infinite right-derivative at $1$. There is however a simple way to show this, without aiming for the first order of $F(\beta,\beta^{\prime})$ as $\beta\downarrow 1$: indeed, using Fatou’s lemma,
$\liminf\limits_{\beta\downarrow 1}\frac{F(\beta,\beta^{\prime})}{\beta-1}\geq\mathbb{E}\mathopen{}\mathclose{{}\left[\liminf\limits_{\beta\downarrow 1}\frac{Q(\beta,\beta^{\prime})}{\beta-1}}\right]=\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{Z(1+\beta^{\prime})}{Z(\beta^{\prime})}}\right],$
@@ 4 Temperature susceptibility
4.1 A first formula for $\kappa_{d}$
@@
$\mathbf{C}_{d}(\beta,\beta+h)=\frac{\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z_{d}(\beta),\log Z_{d}(\beta+h)}\right)}{\sigma(\log Z_{d}(\beta))\,\sigma(\log Z_{d}(\beta+h))}$ (4.1)
and define the susceptibility in temperature, if it exists, as the coefficient $\kappa_{d}(\beta)$ such that
$\mathbf{C}_{d}(\beta,\beta+h)=1-\kappa_{d}(\beta)\,h^{2}+o(h^{2}),\qquad h\rightarrow 0.$ (4.2)
The main technical tool to prove existence of $\kappa_{d}(\beta)$ is the following lemma.
Lemma 4.1.
@@
$\beta\in(1,\infty)\mapsto f_{1}(\beta)\coloneqq\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{d}(\beta^{\prime})\log Z_{d}(\beta)}\right],$
$\beta\in(1,\infty)\mapsto f_{2}(\beta)\coloneqq\mathbb{E}\mathopen{}\mathclose{{}\left[\log^{2}Z_{d}(\beta)}\right],$
$\beta\in(1,\infty)\mapsto f_{3}(\beta)\coloneqq\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{d}(\beta)}\right],$
@@
$Z_{d}(\beta)=\mathrm{e}^{\beta\xi_{1}}\sum_{k\geq 1}\mathrm{e}^{\beta E_{k}},$
@@
$f_{1}(\beta)$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{d}(\beta^{\prime})\log\mathopen{}\mathclose{{}\left(\mathrm{e}^{\beta\xi_{1}}\sum\mathrm{e}^{\beta E_{k}}}\right)}\right]$
$=\beta\,\mathbb{E}\mathopen{}\mathclose{{}\left[\xi_{1}\log Z_{d}(\beta^{\prime})}\right]+\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{d}(\beta^{\prime})\log\sum\mathrm{e}^{\beta E_{k}}}\right].$
@@
$f_{1}^{\prime}(\beta)$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\xi_{1}\log Z_{d}(\beta^{\prime})}\right]+\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{d}(\beta^{\prime})\frac{\sum E_{k}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}}\right],$
@@
$\mathopen{}\mathclose{{}\left|\log Z_{d}(\beta^{\prime})\,\frac{\sum E_{k}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}}\right|$ $\leq\mathopen{}\mathclose{{}\left|\log Z_{d}(\beta^{\prime})}\right|\frac{\sum|E_{k}|\,\mathrm{e}^{\beta_{1}E_{k}}}{\sum\mathrm{e}^{\beta_{1}E_{k}}}\,,$
and those two last terms have finite second moments. The first one because $\log Z_{d}(\beta^{\prime})$ has the same law as $\log Z(\beta)$ up to translation and the second one since
@@
$f_{1}^{\prime\prime}(\beta)$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{d}(\beta^{\prime})\mathopen{}\mathclose{{}\left(\frac{\sum E_{k}^{2}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}-\mathopen{}\mathclose{{}\left(\frac{\sum E_{k}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}}\right)^{2}}\right)}\right],$
@@
$\mathopen{}\mathclose{{}\left|\log Z_{d}(\beta^{\prime})\mathopen{}\mathclose{{}\left(\frac{\sum E_{k}^{2}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}-\mathopen{}\mathclose{{}\left(\frac{\sum E_{k}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}}\right)^{2}}\right)}\right|$ $\leq\mathopen{}\mathclose{{}\left|\log Z_{d}(\beta^{\prime})}\right|\frac{\sum E_{k}^{2}\,\mathrm{e}^{\beta_{1}E_{k}}}{\sum\mathrm{e}^{\beta_{1}E_{k}}}\,,$
@@
$\mathopen{}\mathclose{{}\left|\log\sum\mathrm{e}^{\beta E_{k}}}\right|\frac{\sum E_{k}^{2}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}\leq\mathopen{}\mathclose{{}\left|\log\sum\mathrm{e}^{\beta_{1}E_{k}}}\right|\frac{\sum E_{k}^{2}\,\mathrm{e}^{\beta_{1}E_{k}}}{\sum\mathrm{e}^{\beta_{1}E_{k}}},$
@@
Corollary 4.2.
Assume that $\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}}\right]<\infty$, for every $\beta>1$. Then, for any $\beta>1$, $\kappa_{d}(\beta)$ is well-defined and given by
$\kappa_{d}(\beta)=\frac{1}{2}\mathopen{}\mathclose{{}\left(\frac{\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z_{d}^{\prime}(\beta)}{Z_{d}(\beta)}}\right)}{\operatorname{Var}\,(\log Z_{d}(\beta))}-\mathopen{}\mathclose{{}\left(\frac{\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z_{d}(\beta),\frac{Z_{d}^{\prime}(\beta)}{Z_{d}(\beta)}}\right)}{\operatorname{Var}\,(\log Z_{d}(\beta))}}\right)^{2}}\right).$
@@
We expand $\mathbf{C}_{d}(\beta,\beta+h)$ as $h\to 0$, starting from Equation (4.1) and using Lemma 4.1, writing $Z_{d}$ instead of $Z_{d}(\beta)$ for brevity,
$\mathbf{C}_{d}(\beta,\beta+h)$ $=\frac{\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z_{d},\log Z_{d}+\frac{Z_{d}^{\prime}}{Z_{d}}\,h+\mathopen{}\mathclose{{}\left(\frac{Z_{d}^{\prime\prime}}{Z_{d}}-\mathopen{}\mathclose{{}\left(\frac{Z_{d}^{\prime}}{Z_{d}}}\right)^{2}}\right)\,h^{2}}\right)}{\sigma\mathopen{}\mathclose{{}\left(\log Z_{d}}\right)\,\sigma\mathopen{}\mathclose{{}\left(\log Z_{d}+\frac{Z_{d}^{\prime}}{Z_{d}}\,h+\mathopen{}\mathclose{{}\left(\frac{Z_{d}^{\prime\prime}}{Z_{d}}-\mathopen{}\mathclose{{}\left(\frac{Z_{d}^{\prime}}{Z_{d}}}\right)^{2}}\right)\,h^{2}}\right)}+o\mathopen{}\mathclose{{}\left(h^{2}}\right)$
$=1-\frac{1}{2}\mathopen{}\mathclose{{}\left(\frac{\operatorname{Var}\frac{Z_{d}^{\prime}}{Z_{d}}}{\operatorname{Var}\log Z_{d}}-\mathopen{}\mathclose{{}\left(\frac{\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z_{d},\frac{Z_{d}^{\prime}}{Z_{d}}}\right)}{\operatorname{Var}\log Z_{d}}}\right)^{2}}\right)h^{2}+o\mathopen{}\mathclose{{}\left(h^{2}}\right)\,,$
@@ 4.2 The REM case
$\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z(\beta),\frac{Z^{\prime}(\beta)}{Z(\beta)}}\right)=\frac{1}{2}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\beta}\operatorname{Var}(\log Z_{d}(\beta))=\frac{\pi^{2}}{6}\beta.$
@@ 4.3 Comparison wih the REM in the general decorated case
In this section, we come back to the general decorated case. Our aim is to prove the following proposition which compares the temperature susceptibility $\kappa_{d}(\beta)$ with the one of the REM.
Proposition 4.5.
@@
$\kappa_{d}(\beta)=\kappa(\beta)+\frac{3}{\pi^{2}\beta(\beta+1)}\operatorname{Var}_{\beta}\mathopen{}\mathclose{{}\left(\frac{1}{\beta}\log S_{\beta}-\frac{S^{\prime}_{\beta}}{S_{\beta}}}\right),$
@@
$\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\log Z_{d}(\beta)}\right)=\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\log\Big{(}\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]^{\beta}Z(\beta)\Big{)}}\right)=\operatorname{Var}\,(\log Z(\beta))=\frac{\pi^{2}}{6}(\beta^{2}-1)$
@@
$\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z_{d}(\beta),\frac{Z_{d}^{\prime}(\beta)}{Z_{d}(\beta)}}\right)=\frac{1}{2}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\beta}\operatorname{Var}\,(\log Z_{d}(\beta))=\frac{1}{2}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\beta}\operatorname{Var}\,(\log Z(\beta))=\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z(\beta),\frac{Z^{\prime}(\beta)}{Z(\beta)}}\right).$
@@
$\kappa_{d}(\beta)=\kappa(\beta)+\frac{3}{\pi^{2}\mathopen{}\mathclose{{}\left(\beta^{2}-1}\right)}\mathopen{}\mathclose{{}\left(\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z_{d}^{\prime}(\beta)}{Z_{d}(\beta)}}\right)-\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z^{\prime}(\beta)}{Z(\beta)}}\right)}\right).$ (4.7)
We shall now focus on $\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z_{d}^{\prime}(\beta)}{Z_{d}(\beta)}}\right)$. Notice first that
$Z_{d}^{\prime}(\beta)=\sum_{i,k}\mathopen{}\mathclose{{}\left(\xi_{i}+d_{ik}}\right)\mathrm{e}^{\beta(\xi_{i}+d_{ik})}.$
@@
$\frac{Z_{d}^{\prime}(\beta)}{Z_{d}(\beta)}\;\text{ under }\mathbb{P}\;\overset{({\rm d})}{=}\;\frac{\sum_{i,k}\mathopen{}\mathclose{{}\left(\xi_{i}+c_{\beta}-\frac{1}{\beta}\log S_{\beta,i}+d_{ik}}\right)\mathrm{e}^{\beta(\xi_{i}-\frac{1}{\beta}\log S_{\beta,i}+d_{ik})}}{\sum_{i}\mathrm{e}^{\beta\xi_{i}}}\;\text{ under }\mathbb{P}_{\beta}.$
@@
$\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z_{d}^{\prime}(\beta)}{Z_{d}(\beta)}}\right)$ $=\operatorname{Var}_{\beta}\mathopen{}\mathclose{{}\left(\mathbb{E}_{\beta}\mathopen{}\mathclose{{}\left[\frac{\sum_{i}\xi_{i}\mathrm{e}^{\beta\xi_{i}}}{\sum_{i}\mathrm{e}^{\beta\xi_{i}}}+\frac{\sum_{i}\mathrm{e}^{\beta\xi_{i}}(S_{\beta,i}^{\prime}/S_{\beta,i}-\frac{1}{\beta}\log S_{\beta,i})}{\sum_{i}\mathrm{e}^{\beta\xi_{i}}}\mathrel{}\middle|\mathrel{}\xi}\right]}\right)$
@@
$\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z_{d}^{\prime}(\beta)}{Z_{d}(\beta)}}\right)$ $=\operatorname{Var}_{\beta}\mathopen{}\mathclose{{}\left(\frac{\sum_{i}\xi_{i}\mathrm{e}^{\beta\xi_{i}}}{\sum_{i}\mathrm{e}^{\beta\xi_{i}}}}\right)+\mathbb{E}_{\beta}\mathopen{}\mathclose{{}\left[\frac{\sum_{i}\mathrm{e}^{2\beta\xi_{i}}}{(\sum_{i}\mathrm{e}^{\beta\xi_{i}})^{2}}}\right]\operatorname{Var}_{\beta}\mathopen{}\mathclose{{}\left(\frac{S_{\beta}^{\prime}}{S_{\beta}}-\frac{1}{\beta}\log S_{\beta}}\right)$
@@
$\kappa_{d}(\beta)=\kappa(\beta)+\frac{3}{\pi^{2}\beta(\beta+1)}\operatorname{Var}_{\beta}\mathopen{}\mathclose{{}\left(\frac{1}{\beta}\log S_{\beta}-\frac{S^{\prime}_{\beta}}{S_{\beta}}}\right).$ (4.8)
In the cases which follow, our strategy to find the behavior of $\kappa_{d}(\beta)-\kappa(\beta)$ is to estimate precisely

### v3

@@ 3.4 The REM case
proving that the function $Q(\,\cdot\,,\beta^{\prime})$ is almost surely right differentiable at $1$. On the other hand, Theorem 1.1 implies that $\mathbb{E}[Q(\,\cdot\,,\beta^{\prime})]$ has an infinite right-derivative at $1$. There is however a simple way to show this, without aiming for the first order of $\mathbb{E}\mathopen{}\mathclose{{}\left[Q(\beta,\beta^{\prime})}\right]$ as $\beta\downarrow 1$: indeed, using Fatou’s lemma,
$\liminf\limits_{\beta\downarrow 1}\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[Q(\beta,\beta^{\prime})}\right]}{\beta-1}\geq\mathbb{E}\mathopen{}\mathclose{{}\left[\liminf\limits_{\beta\downarrow 1}\frac{Q(\beta,\beta^{\prime})}{\beta-1}}\right]=\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{Z(1+\beta^{\prime})}{Z(\beta^{\prime})}}\right],$
@@ 4 Temperature susceptibility
4.1 A first formula for $\kappa_{\mathrm{d}}$
@@
$\mathbf{C}_{\mathrm{d}}(\beta,\beta+h)=\frac{\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z_{\mathrm{d}}(\beta),\log Z_{\mathrm{d}}(\beta+h)}\right)}{\sigma(\log Z_{\mathrm{d}}(\beta))\,\sigma(\log Z_{\mathrm{d}}(\beta+h))}$ (4.1)
and define the susceptibility in temperature, if it exists, as the coefficient $\kappa_{\mathrm{d}}(\beta)$ such that
$\mathbf{C}_{\mathrm{d}}(\beta,\beta+h)=1-\kappa_{\mathrm{d}}(\beta)\,h^{2}+o(h^{2}),\qquad h\rightarrow 0.$ (4.2)
The main technical tool to prove existence of $\kappa_{\mathrm{d}}(\beta)$ is the following lemma.
Lemma 4.1 (General decorated case).
@@
$\beta\in(1,\infty)\mapsto f_{1}(\beta)\coloneqq\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{\mathrm{d}}(\beta^{\prime})\log Z_{\mathrm{d}}(\beta)}\right],$
$\beta\in(1,\infty)\mapsto f_{2}(\beta)\coloneqq\mathbb{E}\mathopen{}\mathclose{{}\left[\log^{2}Z_{\mathrm{d}}(\beta)}\right],$
$\beta\in(1,\infty)\mapsto f_{3}(\beta)\coloneqq\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{\mathrm{d}}(\beta)}\right],$
@@
$Z_{\mathrm{d}}(\beta)=\mathrm{e}^{\beta\xi_{1}}\sum_{k\geq 1}\mathrm{e}^{\beta E_{k}},$
@@
$f_{1}(\beta)=\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{\mathrm{d}}(\beta^{\prime})\log\mathopen{}\mathclose{{}\left(\mathrm{e}^{\beta\xi_{1}}\sum\mathrm{e}^{\beta E_{k}}}\right)}\right]=\beta\,\mathbb{E}\mathopen{}\mathclose{{}\left[\xi_{1}\log Z_{\mathrm{d}}(\beta^{\prime})}\right]+\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{\mathrm{d}}(\beta^{\prime})\log\sum\mathrm{e}^{\beta E_{k}}}\right].$
@@
$f_{1}^{\prime}(\beta)=\mathbb{E}\mathopen{}\mathclose{{}\left[\xi_{1}\log Z_{\mathrm{d}}(\beta^{\prime})}\right]+\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{\mathrm{d}}(\beta^{\prime})\frac{\sum E_{k}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}}\right],$
@@
$\mathopen{}\mathclose{{}\left\lvert\log Z_{\mathrm{d}}(\beta^{\prime})\,\frac{\sum E_{k}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}}\right\rvert\leq\mathopen{}\mathclose{{}\left\lvert\log Z_{\mathrm{d}}(\beta^{\prime})}\right\rvert\frac{\sum\,\mathopen{}\mathclose{{}\left\lvert E_{k}}\right\rvert\,\mathrm{e}^{\beta_{1}E_{k}}}{\sum\mathrm{e}^{\beta_{1}E_{k}}}\,,$
and those two last terms have finite second moments. The first one because $\log Z_{\mathrm{d}}(\beta^{\prime})$ has the same law as $\log Z(\beta)$ up to translation and the second one since
@@
$f_{1}^{\prime\prime}(\beta)=\mathbb{E}\mathopen{}\mathclose{{}\left[\log Z_{\mathrm{d}}(\beta^{\prime})\mathopen{}\mathclose{{}\left(\frac{\sum E_{k}^{2}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}-\mathopen{}\mathclose{{}\left(\frac{\sum E_{k}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}}\right)^{2}}\right)}\right],$
@@
$\mathopen{}\mathclose{{}\left\lvert\log Z_{\mathrm{d}}(\beta^{\prime})\mathopen{}\mathclose{{}\left(\frac{\sum E_{k}^{2}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}-\mathopen{}\mathclose{{}\left(\frac{\sum E_{k}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}}\right)^{2}}\right)}\right\rvert\leq\mathopen{}\mathclose{{}\left\lvert\log Z_{\mathrm{d}}(\beta^{\prime})}\right\rvert\frac{\sum E_{k}^{2}\,\mathrm{e}^{\beta_{1}E_{k}}}{\sum\mathrm{e}^{\beta_{1}E_{k}}}\,,$
@@
$\mathopen{}\mathclose{{}\left\lvert\log\sum\mathrm{e}^{\beta E_{k}}}\right\rvert\,\frac{\sum E_{k}^{2}\,\mathrm{e}^{\beta E_{k}}}{\sum\mathrm{e}^{\beta E_{k}}}\leq\mathopen{}\mathclose{{}\left\lvert\log\sum\mathrm{e}^{\beta_{1}E_{k}}}\right\rvert\,\frac{\sum E_{k}^{2}\,\mathrm{e}^{\beta_{1}E_{k}}}{\sum\mathrm{e}^{\beta_{1}E_{k}}},$
@@
Corollary 4.2 (General decorated case).
Assume that $\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}}\right]<\infty$, for every $\beta>1$. Then, for any $\beta>1$, $\kappa_{\mathrm{d}}(\beta)$ is well-defined and given by
$\kappa_{\mathrm{d}}(\beta)=\frac{1}{2}\mathopen{}\mathclose{{}\left(\frac{\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z_{\mathrm{d}}^{\prime}(\beta)}{Z_{\mathrm{d}}(\beta)}}\right)}{\operatorname{Var}\,(\log Z_{\mathrm{d}}(\beta))}-\mathopen{}\mathclose{{}\left(\frac{\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z_{\mathrm{d}}(\beta),\frac{Z_{\mathrm{d}}^{\prime}(\beta)}{Z_{\mathrm{d}}(\beta)}}\right)}{\operatorname{Var}\,(\log Z_{\mathrm{d}}(\beta))}}\right)^{2}}\right).$
@@
We expand $\mathbf{C}_{\mathrm{d}}(\beta,\beta+h)$ as $h\to 0$, starting from Equation (4.1) and using Lemma 4.1, writing $Z_{\mathrm{d}}$ instead of $Z_{\mathrm{d}}(\beta)$ for brevity,
$\mathbf{C}_{\mathrm{d}}(\beta,\beta+h)$ $=\frac{\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z_{\mathrm{d}},\log Z_{\mathrm{d}}+\frac{Z_{\mathrm{d}}^{\prime}}{Z_{\mathrm{d}}}\,h+\mathopen{}\mathclose{{}\left(\frac{Z_{\mathrm{d}}^{\prime\prime}}{Z_{\mathrm{d}}}-\mathopen{}\mathclose{{}\left(\frac{Z_{\mathrm{d}}^{\prime}}{Z_{\mathrm{d}}}}\right)^{2}}\right)\,h^{2}}\right)}{\sigma\mathopen{}\mathclose{{}\left(\log Z_{\mathrm{d}}}\right)\,\sigma\mathopen{}\mathclose{{}\left(\log Z_{\mathrm{d}}+\frac{Z_{\mathrm{d}}^{\prime}}{Z_{\mathrm{d}}}\,h+\mathopen{}\mathclose{{}\left(\frac{Z_{\mathrm{d}}^{\prime\prime}}{Z_{\mathrm{d}}}-\mathopen{}\mathclose{{}\left(\frac{Z_{\mathrm{d}}^{\prime}}{Z_{\mathrm{d}}}}\right)^{2}}\right)\,h^{2}}\right)}+o\mathopen{}\mathclose{{}\left(h^{2}}\right)$
$=1-\frac{1}{2}\mathopen{}\mathclose{{}\left(\frac{\operatorname{Var}\frac{Z_{\mathrm{d}}^{\prime}}{Z_{\mathrm{d}}}}{\operatorname{Var}\log Z_{\mathrm{d}}}-\mathopen{}\mathclose{{}\left(\frac{\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z_{\mathrm{d}},\frac{Z_{\mathrm{d}}^{\prime}}{Z_{\mathrm{d}}}}\right)}{\operatorname{Var}\log Z_{\mathrm{d}}}}\right)^{2}}\right)h^{2}+o\mathopen{}\mathclose{{}\left(h^{2}}\right)\,,$
@@ 4.2 The REM case
$\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z(\beta),\frac{Z^{\prime}(\beta)}{Z(\beta)}}\right)=\frac{1}{2}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\beta}\operatorname{Var}(\log Z(\beta))=\frac{\pi^{2}}{6}\beta.$
@@ 4.3 Comparison wih the REM in the general decorated case
In this section, we come back to the general decorated case. Our aim is to prove the following proposition which compares the temperature susceptibility $\kappa_{\mathrm{d}}(\beta)$ with the one of the REM.
Proposition 4.5 (General decorated case).
@@
$\kappa_{\mathrm{d}}(\beta)=\kappa(\beta)+\frac{3}{\pi^{2}\beta(\beta+1)}\operatorname{Var}_{\beta}\mathopen{}\mathclose{{}\left(\frac{1}{\beta}\log S_{\beta}-\frac{S^{\prime}_{\beta}}{S_{\beta}}}\right),$
@@
$\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\log Z_{\mathrm{d}}(\beta)}\right)=\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\log\Big{(}\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]^{\beta}Z(\beta)\Big{)}}\right)=\operatorname{Var}\,(\log Z(\beta))=\frac{\pi^{2}}{6}(\beta^{2}-1)$
@@
$\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z_{\mathrm{d}}(\beta),\frac{Z_{\mathrm{d}}^{\prime}(\beta)}{Z_{\mathrm{d}}(\beta)}}\right)=\frac{1}{2}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\beta}\operatorname{Var}\,(\log Z_{\mathrm{d}}(\beta))=\frac{1}{2}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\beta}\operatorname{Var}\,(\log Z(\beta))=\operatorname{Cov}\mathopen{}\mathclose{{}\left(\log Z(\beta),\frac{Z^{\prime}(\beta)}{Z(\beta)}}\right).$
@@
$\kappa_{\mathrm{d}}(\beta)=\kappa(\beta)+\frac{3}{\pi^{2}\mathopen{}\mathclose{{}\left(\beta^{2}-1}\right)}\mathopen{}\mathclose{{}\left(\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z_{\mathrm{d}}^{\prime}(\beta)}{Z_{\mathrm{d}}(\beta)}}\right)-\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z^{\prime}(\beta)}{Z(\beta)}}\right)}\right).$ (4.7)
We shall now focus on $\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z_{\mathrm{d}}^{\prime}(\beta)}{Z_{\mathrm{d}}(\beta)}}\right)$. Notice first that
$Z_{\mathrm{d}}^{\prime}(\beta)=\sum_{i,k}\mathopen{}\mathclose{{}\left(\xi_{i}+d_{ik}}\right)\mathrm{e}^{\beta(\xi_{i}+d_{ik})}.$
@@
$\frac{Z_{\mathrm{d}}^{\prime}(\beta)}{Z_{\mathrm{d}}(\beta)}\;\text{ under }\mathbb{P}\;\overset{({\rm d})}{=}\;\frac{\sum_{i,k}\mathopen{}\mathclose{{}\left(\xi_{i}+c_{\beta}-\frac{1}{\beta}\log S_{\beta,i}+d_{ik}}\right)\mathrm{e}^{\beta(\xi_{i}-\frac{1}{\beta}\log S_{\beta,i}+d_{ik})}}{\sum_{i}\mathrm{e}^{\beta\xi_{i}}}\;\text{ under }\mathbb{P}_{\beta}.$
@@
$\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z_{\mathrm{d}}^{\prime}(\beta)}{Z_{\mathrm{d}}(\beta)}}\right)$ $=\operatorname{Var}_{\beta}\mathopen{}\mathclose{{}\left(\mathbb{E}_{\beta}\mathopen{}\mathclose{{}\left[\frac{\sum_{i}\xi_{i}\mathrm{e}^{\beta\xi_{i}}}{\sum_{i}\mathrm{e}^{\beta\xi_{i}}}+\frac{\sum_{i}\mathrm{e}^{\beta\xi_{i}}(S_{\beta,i}^{\prime}/S_{\beta,i}-\frac{1}{\beta}\log S_{\beta,i})}{\sum_{i}\mathrm{e}^{\beta\xi_{i}}}\mathrel{}\middle|\mathrel{}\xi}\right]}\right)$
@@
$\operatorname{Var}\,\mathopen{}\mathclose{{}\left(\frac{Z_{\mathrm{d}}^{\prime}(\beta)}{Z_{\mathrm{d}}(\beta)}}\right)$ $=\operatorname{Var}_{\beta}\mathopen{}\mathclose{{}\left(\frac{\sum_{i}\xi_{i}\mathrm{e}^{\beta\xi_{i}}}{\sum_{i}\mathrm{e}^{\beta\xi_{i}}}}\right)+\mathbb{E}_{\beta}\mathopen{}\mathclose{{}\left[\frac{\sum_{i}\mathrm{e}^{2\beta\xi_{i}}}{(\sum_{i}\mathrm{e}^{\beta\xi_{i}})^{2}}}\right]\operatorname{Var}_{\beta}\mathopen{}\mathclose{{}\left(\frac{S_{\beta}^{\prime}}{S_{\beta}}-\frac{1}{\beta}\log S_{\beta}}\right)$
@@
$\kappa_{\mathrm{d}}(\beta)=\kappa(\beta)+\frac{3}{\pi^{2}\beta(\beta+1)}\operatorname{Var}_{\beta}\mathopen{}\mathclose{{}\left(\frac{1}{\beta}\log S_{\beta}-\frac{S^{\prime}_{\beta}}{S_{\beta}}}\right).$ (4.8)
In the cases which follow, our strategy to find the behavior of $\kappa_{\mathrm{d}}(\beta)-\kappa(\beta)$ is to estimate precisely